OpenStudy (anonymous):

Find the directional derivative of f(x,y)=e^(xsiny) at the point P(1,0) in the direction of Q(-2,4)

OpenStudy (turingtest):

Partial w/ respect to x: \[\sin y [e^(xsiny)]\] evaluated at (1,0)=0 w/ repect to y: \[x \cos y[e^(x \sin y)]\] evaluated at (1,0)=1 \[D _{Q}f(1,0)=af _{x}+b f _{y}=(-2)(0)+(4)(1)=4\] where a and b are the components of Q

OpenStudy (anonymous):

thank you :)